determine the number of atoms in 1.75 mol Al
1
Expert's answer
2020-09-30T04:18:57-0400

"1.75 moles Al \times \frac{6.022\times 10^{23} atoms (Al)}{1 mole(Al)}= 1.05\times 10^{24} atoms (Al)"